what is the power output of a uranium reactor if it takes 30 days to use up 2kg of fuel and each fission gives 185mev of usable energy

Grade:

2 Answers

AskIITiansExpert Kartik-IITMadras
13 years ago


No of mole of uranium 235 in 2 kg = 2*1000/235= 8.51
Each fission gives 185 Mev which is  185 × 1.60× 10^-13 joules E = 296 × 10^-13 J
One mole gives 6.023×10^23 × E = 1.78×10^13 J

Therefore total energy released with 2 kg U-235= 8.51×1.78×10^13 = 1.517×10^14 J.

Power = energy/ time.

Since this energy is released over 30 days time = 30×24×60×60= 2592000s
Hence power output = 1.517×10^14/2.592×10^6 W
                              =5.85×10^7 W

Abhinandan
7 years ago
No of mole of uranium 235 in 2 kg = 2*1000/235= 8.51Each fission gives 185 Mev which is 185 × 1.60× 10^-13 joules E = 296 × 10^-13 JOne mole gives 6.023×10^23 × E = 1.78×10^13 JTherefore total energy released with 2 kg U-235= 8.51×1.78×10^13 = 1.517×10^14 J. Power = energy/ time.Since this energy is released over 30 days time = 30×24×60×60= 2592000sHence power output = 1.517×10^14/2.592×10^6 W =5.85×10^7 W